# Establishing Task Scaling Laws via Compute-Efficient Model Ladders

### v1

@@ Abstract
We develop task scaling laws and model ladders to predict the individual task performance of pretrained language models (LMs) in the overtrained setting. Standard power laws for language modeling loss cannot accurately model task performance. Therefore, we leverage a two-step prediction approach: first use model and data size to predict a task-specific loss, and then use this task loss to predict task performance. We train a set of small-scale “ladder” models, collect data points to fit the parameterized functions of the two prediction steps, and make predictions for two target models: a 7B model trained to 4T tokens and a 13B model trained to 5T tokens. Training the ladder models only costs 1% of the compute used for the target models. On four multiple-choice tasks written in ranked classification format, we can predict the accuracy of both target models within 2 points of absolute error. We have higher prediction error on four other tasks (average absolute error 6.9) and find that these are often tasks with higher variance in task metrics. We also find that using less compute to train fewer ladder models tends to deteriorate predictions. Finally, we empirically show that our design choices and the two-step approach lead to superior performance in establishing scaling laws.
@@ 1 Introduction
Language models (LMs) are expensive to pretrain. Given time and compute constraints, and given the variety of decisions that go into building an LM (e.g., picking a “good” mixture of pretraining data, making optimal model architecture choices), being able to predict the task performance of a large model before actually training it enables more efficient resource allocation and a wider range of experimentation. However, it is challenging to extrapolate the task performance of a large training run from models trained at smaller scales (Hu et al., 2023). Previous work has shown preliminary results on predicting the average performance over many tasks (Gadre et al., 2024), or for a single selected task (ARC-Challenge) (Dubey et al., 2024), using at least 5% of the compute required to train the target models. It remains an open problem to develop a general and compute-efficient method to predict the performance of LMs on individual tasks.
In this paper, we tackle the challenge of predicting individual task performance of LMs as a function of model size and training data size. In particular, we predict the performance of 7B and 13B models using 1% of the compute required to train them. In this work, we focus on multiple-choice tasks, and aim to predict the task accuracy where problems are written in the ranked classification (RC) format. We employ a two-step approach to (1) use the number of model parameters and training tokens to predict a task-specific loss, which is a bits-per-byte loss on the correct answer given the question, then (2) use this task loss to predict accuracy. The closest existing work (Dubey et al., 2024) only examined compute-optimal models; our approach also works for the overtrained regime, which is important because most recent LMs are heavily overtrained (Dubey et al., 2024; Li et al., 2024; Bai et al., 2023).
To fit the parameters of the two functions used to make our predictions, we measure the task loss and accuracy of a variety of small models, varying model parameters between 190M and 1.3B non-embedding parameters, and data sizes between 1x and 10x of the Chinchilla-optimal data size for each model (Hoffmann et al., 2022). Training these ladder models altogether costs only 1% the combined compute of the two target model that we predict for: a 7B model trained on 4T tokens, and a 13B model trained on 5T tokens. 111We will release our code and experiments on Github.
We apply our method on 8 selected tasks from OLMES (Gu et al., 2024) (e.g., MMLU, HellaSwag, ARC-Challenge). On 4 of these tasks, our predictions are within an absolute error of 2 points for the 7B-4T and 13B-5T models. Error on the other 4 tasks is higher, with an average absolute error of 6.9 points, and our analyses show that the metrics of these tasks have relatively high variance between neighboring model checkpoints.
In §5, we conduct several analyses. We quantify the difficulty of the prediction task itself for the different target models by measuring the variance in the task loss and accuracy of the final few checkpoints (the higher the variance in the prediction target, the harder the task) (§5.1). We show that using less compute to train fewer ladder models results in worse predictions (§5.2).
In §6, we study the impact of various design choices in our prediction method, such as using compute-FLOPs instead of as the input (§6.1), using different intermediate features such as task cross-entropy loss and general language modeling loss (§6.2 and §6.3), and bypassing the intermediate feature altogether by making predictions in a single step (§6.4). We find that some tasks benefit from these alternative choices, but for the tasks that we care the most about (e.g., MMLU and HellaSwag), our two-step approach using task loss works better.
@@ 2 Setup
We aim to predict the task performance of LMs with an arbitrary training scale – the combination of model size () and number of training tokens (). Since many recent LMs are heavily overtrained (Dubey et al., 2024; Li et al., 2024; Bai et al., 2023), we do not constrain and to stay close to the compute-optimal regime (Hoffmann et al., 2022). In particular, we make and validate our predictions on the OLMo 2 models after stage 1 pretraining and before stage 2 annealing (Ai2, 2024); these include a 7B model trained with 4T tokens (“7B-4T”) and a 13B model trained with 5T tokens (“13B-5T”), both trained from scratch with the same data mixture. We refer to the 7B-4T and 13B-5T models as target models.
We consider multiple-choice tasks and compute task accuracy in the ranked classification (RC) format (§2.2). The relationship between a model’s accuracy on a given task and its training scale is complex, so we use an intermediate task loss as a bridge. This task loss is defined in §2.2. We predict a given model’s accuracy on a task by chaining two steps: step 1, using model size and training tokens to predict the task loss; step 2, using the task loss to predict accuracy. We define parametric functions for each step (§3.1, §3.2); we fit these parametric functions by training a variety of small models and measuring their task loss and accuracy (§2.1). For these models, we vary and : model sizes between 190M and 1.3B parameters, trained to between 1x and 10x of the Chinchilla-optimal data size (but otherwise keeping the architecture, training data, etc. fixed).
@@ 2.1 Ladder models
To predict the task loss and accuracy of large models, we extrapolate from data points collected from training many small-scale models (“ladder models”). These ladder models have the same architecture and are trained with the same data mix as the target models. The ladder models need to span a relatively wide range of model size and training data size, while collectively costing a very small fraction of compute used for training the large target models.
We use four different model sizes { 190M, 370M, 760M, 1.3B } (considering only non-embedding parameters) by varying the width and depth of the transformer. For each model size, we train it to different multiples of the “Chinchilla optimal” number of tokens. We use as the Chinchilla optimal setting (Hoffmann et al., 2022) (denoted “1xC”), and train each model with number of tokens { 1xC, 2xC, 5xC, 10xC }. In total, we train ladder models. We save intermediate checkpoints every 200 steps for 1xC and 2xC runs, every 500 steps for 5xC runs, and every 1000 steps for 10xC runs. Table 1 lists the configurations of each model size.
@@
The target 7B-4T and 13B-5T models are trained with a cosine LR schedule with linear warmup, where the LR decays to 10% of the peak LR over a horizon of 5T tokens. We match this in the ladder models, where the decay horizon is adjusted to match the training data size of each model. Unfortunately, the OLMo 2 7B model was only trained to 3.9T tokens in stage 1 pretraining. Prior work has shown that LR decay has a big impact on the performance of pretrained LMs. To account for the incomplete training, we train this model with an additional 50B tokens where we linearly decay the LR down to 10%. Our target 7B-4T model is thus trained on a total of 3.95T tokens.
@@
Using to estimate the compute (Kaplan et al., 2020), the total amount of compute used for training the ladder models is FLOPs. This is merely 3.2% of that used for training the 7B-4T model ( FLOPs), 1.3% of the 13B-5T model ( FLOPs), and less than 1.0% of the two target models combined.
2.2 Task loss and accuracy
In this work we focus on multiple-choice tasks. Here, we describe how we format multiple-choice problems to compute task loss and accuracy.
@@ 2.2 Task loss and accuracy
For a multiple-choice problem, we define its task loss as the negative log-likelihood of the correct answer sequence, divided by its length in bytes. This is also known as the bits-per-byte (bpb) metric. Table 2 shows the problem formatting for computing task loss on an example. We use bpb instead of normalizing by number of tokens because bpb reduces the impact of tokenizer.222We anticipate this to be similar to the “Normalized NLL per Char” used in Dubey et al. (2024).
In §6 we also study the effectiveness of alternative choices to task loss, like considering both correct and incorrect answers in the task loss, and using a general language modeling loss. We found that some tasks may benefit from these alternative choices, but task loss works best for predicting task accuracy on MMLU, which is the task we care most about.
@@
For task accuracy, we use the two commonly used formats: ranked classification (RC) and multiple-choice (MC). (Note that RC and MC are different formats to pose multiple-choice problems.) In RC, the predicted answer is the one with the minimum bpb loss. In MC, all the answer choices are included in the prompt, and the predicted answer is the answer code (e.g., A, B, C, etc.) with the smallest loss. See Gu et al. (2024) for detailed discussion on the RC vs. MC formats. In this paper we focus on the RC format.333RC reliably measures progress for models across a wide range of scales, whereas MC capability does not emerge until the model has several billion parameters.
@@
Following Gadre et al. (2024), we evaluate the goodness of prediction by computing its relative error against the actual task accuracy of the target model. The relative error is defined as
@@ 2.3 Task selection
We include the following 8 tasks from the OLMES evaluation suite (Gu et al., 2024): MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), ARC-Challenge (Clark et al., 2018), ARC-Easy (Clark et al., 2018), PIQA (Bisk et al., 2020), CommonsenseQA (Talmor et al., 2019), Social IQa (Sap et al., 2019), and OpenBookQA (Mihaylov et al., 2018). We exclude two tasks from the OLMES suite, BoolQ (Clark et al., 2019) and Winogrande (Sakaguchi et al., 2020), as we do not observe monotonic scaling of task performance on the ladder models. §5.1 has further discussion on task predictability with the ladder.
We use the test set whenever possible, and fall back to the validation set when the test set is not fully available. For all tasks, we use the 5-shot setting as provided in OLMES. Table 3 summarizes the tasks and statistics.
| Task | Split | Examples |
| MMLU | test | 14,042 |
| HellaSwag | val | 10,042 |
| ARC-Challenge | test | 1,172 |
| ARC-Easy | test | 2,376 |
| PIQA | val | 1,838 |
| CommonsenseQA | val | 1,221 |
| Social IQa | val | 1,954 |
| OpenBookQA | test | 500 |
@@ 3 Method
We break down the problem of task accuracy prediction into two steps: first predicting the task loss (taken as the intermediate feature), and then using that to predict the task accuracy. §6 discusses various alternative choices for the intermediate feature.
3.1 Step 1: Use to predict task loss
As proposed by Hoffmann et al. (2022) and followed by others (Muennighoff et al., 2023; Gadre et al., 2024; Zhang et al., 2024), the language modeling loss of a model on a held-out eval set can be modeled as a power function with respect to and :
@@ 3.1 Step 1: Use to predict task loss
We postulate that the same functional form applies to the task loss. We will later validate this assumption by looking at the goodness of function fitting. (In §6.1 we ablate with using compute-FLOPs as input variable instead of , and in §6.2 and §6.3 we ablate with using alternative losses as the intermediate feature.)
@@
We take the loss value of the final checkpoint of each ladder model, forming a dataset (where ), and use this data to fit the parameters of Equation 1. We fit a separate set of parameters for each task. For parameter fitting, we follow the standard practice set in Hoffmann et al. (2022). We minimize the Huber loss between the logarithm of predicted and actual loss: , where . We optimize and in log space – we apply transformation and optimize . We use the L-BFGS-B optimizer as implemented in scipy.minimize(), with constraints ; with these constraints, the function is convex, and thus L-BFGS-B is guaranteed to converge.444To test if this function is convex, we calculate the Hessian matrix of second derivatives, and confirm that it is positive semi-definite everywhere.
To improve function fitting, we take the average task loss over the last 5 checkpoints of each ladder model. This reduces the noisy variance between model checkpoints.
@@
The parameterized power function can fit the ladder models well, with relative fitting error ranging from 0.2% to 1.2% on all tasks. When predicting task loss of target models, across all tasks we get an average relative error of 5.2% for 7B-4T, and 6.6% for 13B-5T. We present results on individual tasks in §4, and ablate the design choices of step 1 in §6.
3.2 Step 2: Use task loss to predict accuracy
Following Dubey et al. (2024), we model the mapping from task loss to accuracy with a sigmoidal function:
@@ 3.2 Step 2: Use task loss to predict accuracy
The choice of a sigmoidal functional form is motivated as follows: a weak model has high task loss and random task accuracy, and a strong model has low task loss and high task accuracy that saturates at 100%. We observed (as shown in Figure 4) that the points collected from different ladder models do tend to fall on a shared sigmoidal curve; this also applies to intermediate checkpoints of these models in addition to final checkpoints.
@@
To fit this function, we use data points from both final and intermediate checkpoints of the ladder models. The task loss and accuracy value of all intermediate and final checkpoints of all ladder models form a dataset (where ), which we use to fit the parameters of Equation 2. We fit a separate set of parameters for each downstream task. We minimize the L2 loss between the predicted and actual accuracy: . We use non-linear least squares implemented by scipy.optimize.curve_fit() to fit this equation, as sigmoid functions are not convex.
The variation from checkpoint to checkpoint can be high. To smoothen the noise, we apply a moving average on the task loss and task accuracy over all checkpoints of each training run, with a window size of 5. We also discard the checkpoints from the first 10% of each training run as these are quite noisy, and add an extra data point to the training pairs. This helps avoid the cases when the fitting function degenerates for very noisy data, and the ladder models are too small to start doing well on certain tasks.
@@
The parameterized sigmoidal function can fit the ladder models well, with relative fitting error ranging from 0.4% to 2.6% on all tasks. When predicting task accuracy of target models with their actual task loss, across all tasks we get an average relative error of 3.7% for 7B-4T, and 3.5% for 13B-5T. We present results on individual tasks in §4, and ablate the design choices of step 1 in §6.

### v2

@@ Abstract
We develop task scaling laws and model ladders to predict the individual task performance of pretrained language models (LMs) in the overtrained setting. Standard power laws for language modeling loss cannot accurately model task performance. Therefore, we leverage a two-step prediction approach: (1) use model and data size to predict an intermediate loss, then (2) use it to predict task performance. We train a set of small-scale “ladder” models, collect data points to fit the parameterized functions of the two prediction steps, and make predictions for two target models: a 7B model trained to 4T tokens and a 13B model trained to 5T tokens. Training the ladder models only costs 1% of the compute used for the target models. On four multiple-choice tasks formatted as ranked classification, we can predict the accuracy of both target models within 2 points of absolute error. We find that tasks with higher prediction error also have higher variance in the metrics over model checkpoints. We also contrast multiple design choices for predicting accuracy, and present recommendations for extending our method to new models and tasks.
@@ 1 Introduction
Language models (LMs) are expensive to train. Building a good LM requires a wide range of experimentation over data and architecture choices. Scaling laws, which use the performance of smaller models to predict the performance of a larger model without actually training it, enable more efficient resource allocation for such experimentation. Further, modern LMs are compared based on their performance on downstream tasks, rather than perplexity. Certain downstream tasks (e.g., HellaSwag) are also often used as development benchmarks for creating training recipes and good data mixtures for training large LMs (OLMo et al., 2024; Li et al., 2024). Scaling laws for downstream tasks are, therefore, important for LM pretraining.
Prior work has shown results on predicting the average top-1 error over many tasks (as opposed to predicting task accuracy directly) (Gadre et al., 2024), or the accuracy for only one specific task (ARC-Challenge) (Dubey et al., 2024), using at least 5% of the compute required to train the target models. Additionally, these require developing a testbed of many small models to first find compute-optimal models, which adds significant compute costs. Predicting LM performance on a range of individual downstream tasks in a computationally efficient way (e.g., without finding compute-optimal models) remains an open problem.
Here, we tackle the challenge of predicting individual task performance of LMs as a function of model size and training data size. In particular, we predict the performance of 7B and 13B models. We focus on a range of multiple-choice tasks and predict the task accuracy for problems written in the ranked classification (RC) format. We use a set of small models (190M to 1.3B non-embedding parameters), which we train for varying durations (1x to 10x Chinchilla optimal data size) to develop scaling laws. Training these fixed-size ladder models costs only 1% of the compute of the two target models combined.
We employ a two-step approach to (1) use the number of model parameters and training tokens (the input features) to predict a task-specific loss (an intermediate feature), and then (2) use this task loss to predict accuracy. We experiment with different design choices for the intermediate feature, as well as input features, and apply our method on 8 selected tasks from OLMES (Gu et al., 2024). We quantify the predictability of the tasks based on the variance of target metrics for a representative small model over the last few training checkpoints. Finally, we provide recommendations for developing downstream scaling laws for new models, and selecting a design choice for a given task. From our recommendations, using a task-specific loss as the intermediate feature, we achieve an absolute error of 2 points for our target models on four tasks (MMLU, HellaSwag, PiQA, SocialIQA), and an average absolute error of 4 points across both target models and all tasks.
@@ 2 Setup
We aim to predict the task performance of LMs with an arbitrary training scale – the combination of model size () and number of training tokens (). As recent LMs are overtrained (Dubey et al., 2024; Li et al., 2024; Bai et al., 2023), we do not constrain and to stay close to the compute-optimal regime (Hoffmann et al., 2022). We validate our predictions on the OLMo 2 models after stage 1 pretraining and before stage 2 annealing (OLMo et al., 2024); a 7B model trained to 4T tokens (“7B-4T”) and a 13B model trained to 5T tokens (“13B-5T”), both trained from scratch with the same data mixture. These are target models.
@@ 2.1 Ladder models
To predict the performance of large models, we extrapolate from data points collected from training many small models (“ladder models”). These models have the same architecture and are trained with the same data mix as the target models. This is useful for guiding pretraining development, as data mixtures and modeling decisions can be tested in a controlled setting. The models span a relatively wide range of model size and training data size, while collectively costing a small fraction of compute used to train the target models.
We define four model sizes { 190M, 370M, 760M, 1.3B } (considering only non-embedding parameters) by varying the width and depth of the transformer. We use as the Chinchilla optimal setting (Hoffmann et al., 2022) (denoted “1xC”), and train each model with number of tokens { 1xC, 2xC, 5xC, 10xC }. In total, we train models. We save intermediate checkpoints every 200 steps for 1xC and 2xC runs, every 500 steps for 5xC runs, and every 1000 steps for 10xC runs. Table 1 lists the configurations of each model size.
@@
The target models are trained with a cosine LR schedule with linear warmup, where the LR decays to 10% of the peak LR over a horizon of 5T tokens. We match this in the ladder models, where the decay horizon is adjusted to match the training data size of each model. Unfortunately, the OLMo 2 7B model was only trained to 3.9T tokens in stage 1 pretraining, whereas the cosine scheduler was set to 5T tokens. To account for this incomplete training, we train this model with an additional 50B tokens where we linearly decay the LR down to 10%. Our target 7B-4T model is thus trained on a total of 3.95T tokens, where the linear decay instead of the slower cosine decay allows us to use fewer tokens.
@@
Using to estimate the compute (Kaplan et al., 2020), the total amount of compute used for training the ladder models is FLOPs. This is only 3.2% of that used for training the 7B-4T model ( FLOPs), 1.3% of the 13B-5T model ( FLOPs), and less than 1.0% of both target models combined.
2.2 Intermediate features and accuracy
@@
We define task loss as the negative log-likelihood of the correct answer sequence, divided by its length in bytes, also known as the bits-per-byte (bpb) metric. Table 2 shows the problem formatting for computing task loss on an example. We use bpb instead of normalizing by number of tokens to reduce the impact of the tokenizer.111similar to “normalized NLL per char” (Dubey et al., 2024).
TaskCE.
Task cross-entropy loss which accounts for incorrect answers (§C.2).
LM loss.
Standard language modeling loss on the C4-en validation set (Raffel et al., 2019).
@@
Ranked classification (RC) and multiple-choice (MC) are two different formats to pose multiple-choice problems.) In RC, the predicted answer is the one with the minimum task loss. In MC, all the answer choices are included in the prompt, and the predicted answer is the answer code (e.g., A, B, C, etc.) with the smallest loss (Gu et al., 2024). Here, we focus on the RC format.222RC reliably measures progress for models across a wide range of scales, whereas MC capability does not emerge until the model has several billion parameters.
@@
Following Gadre et al. (2024), we compute the relative error for the task accuracy to evaluate the goodness of prediction, defined as
@@ 2.3 Task selection
We include the following 8 tasks from the OLMES evaluation suite (Gu et al., 2024): MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), ARC-Challenge (Clark et al., 2018), ARC-Easy (Clark et al., 2018), PIQA (Bisk et al., 2020), CommonsenseQA (Talmor et al., 2019), Social IQa (Sap et al., 2019), and OpenBookQA (Mihaylov et al., 2018). We exclude BoolQ (Clark et al., 2019) and Winogrande (Sakaguchi et al., 2020), as the task loss and accuracy are noisier for these tasks. §5 discusses task predictability further. We use the test set where possible, and fall back to the validation set otherwise. For all tasks, we use the 5-shot setting provided in OLMES. Figure 2 shows the task statistics.
| 7B-4T | 13B-5T | |||||||||
| Task | Split | Examples | Pred | Actual | Error | %Error | Pred | Actual | Error | %Error |
| MMLU | test | 14,042 | 48.4 | 49.0 | 0.6 | 1.3% | 51.3 | 51.6 | 0.3 | 0.7% |
| HellaSwag | val | 10,042 | 82.5 | 81.3 | 1.2 | 1.4% | 85.3 | 83.2 | 2.1 | 2.5% |
| ARC-Challenge | test | 1,172 | 51.5 | 61.9 | 10.4 | 16.9% | 52.7 | 63.8 | 11.1 | 17.5% |
| ARC-Easy | test | 2,376 | 76.6 | 84.6 | 8.0 | 9.4% | 77.2 | 87.2 | 10.0 | 11.4% |
| PIQA | val | 1,838 | 81.2 | 82.0 | 0.8 | 1.0% | 82.1 | 83.0 | 0.9 | 1.1% |
| CommonsenseQA | val | 1,221 | 75.7 | 72.6 | 3.1 | 4.2% | 77.6 | 74.1 | 3.5 | 4.7% |
| Social IQa | val | 1,954 | 58.7 | 59.9 | 1.2 | 2.0% | 59.9 | 61.6 | 1.7 | 2.7% |
| OpenBookQA | test | 500 | 44.2 | 49.4 | 5.2 | 10.6% | 44.9 | 48.6 | 3.7 | 7.8% |
| Average | – | – | – | – | 3.8 | 5.9% | – | – | 4.2 | 6.1% |
@@ 3 Method
We break down task accuracy prediction into two steps: 1) predicting the intermediate feature (we use task loss to illustrate this), and 2) using it to predict the task accuracy.
3.1 Step 1: Use to predict the intermediate feature
We consider three intermediate features: task loss, taskCE loss, and LM loss (defined in §2).
As proposed by Hoffmann et al. (2022) and followed by others (Muennighoff et al., 2023; Gadre et al., 2024; Zhang et al., 2024), the LM loss of a model on a held-out eval set can be modeled as a power function with respect to and :
@@ 3.1 Step 1: Use to predict the intermediate feature
We postulate that the same functional form applies to these losses. We validate this assumption by looking at the goodness of function fitting.
@@
We take the loss value of the final checkpoint of each ladder model (where ), and use this data to fit the parameters of Equation 1. We fit a separate set of parameters for each task. Following Hoffmann et al. (2022), we minimize the Huber loss between the logarithm of predicted and actual loss: , where . We optimize and in log space – we apply transformation and optimize . We use the L-BFGS-B optimizer as implemented in scipy.minimize(), with constraints ; with these constraints, the function is convex, and thus L-BFGS-B is guaranteed to converge.333To test if this function is convex, we confirm that the Hessian matrix of second derivatives is positive semi-definite everywhere.
We take the average loss over the last 5 checkpoints of each ladder model to reduce noise.
@@
Figure 3 shows the function fitting of step 1 on the ladder models, and the prediction of task loss as the intermediate feature for the target models. The power function gives an average relative fitting error ranging from 0.2% to 1.2%, which suggests that Equation 1 is a good functional form to describe task losses. We have a relative prediction error within 3% on MMLU, HellaSwag, PIQA, and OpenBookQA. The method underestimates the loss for CSQA, and overestimates on ARC-Challenge, ARC-Easy, and Social IQa. On target models, across all tasks we get an average relative error of 5.2% for 7B-4T, and 6.6% for 13B-5T. Results for the other intermediate features are discussed in §C.2 and §C.3.
In §C.1, we use compute-FLOPs as input variable instead of , and observe higher fitting errors for all tasks. We also note that FLOPs cannot distinguish between compute-optimal and overtrained models.
3.2 Step 2: Use intermediate feature to predict accuracy
Following Dubey et al. (2024), we model the mapping from the intermediate loss to accuracy with a sigmoidal function:
@@ 3.2 Step 2: Use intermediate feature to predict accuracy
The choice of a sigmoidal functional form is motivated as follows: a weak model has high task loss and random task accuracy, and a strong model has low task loss and high task accuracy that saturates at 100%. We observe (as shown in Figure 4) that the points collected from different ladder models tend to fall on a shared sigmoidal curve; this applies to both intermediate and final model checkpoints.
@@
We use the intermediate loss and accuracy values from both final and intermediate checkpoints of the ladder models (where ) to fit the parameters of Equation 2. We fit a separate set of parameters for each downstream task. We minimize the L2 loss between the predicted and actual accuracy: . We use non-linear least squares implemented by scipy.optimize.curve_fit() to fit this equation, as sigmoid functions are not convex.
The variation from checkpoint to checkpoint can be high. To smoothen the noise, we apply a moving average on the intermediate loss and task accuracy over all checkpoints of each training run, with a window size of 5. We also discard the checkpoints from the first 10% of each training run as these are quite noisy, and add an extra data point to the training pairs. This helps avoid the cases when the fitting function degenerates for very noisy data, and the ladder models are too small to do well on certain tasks.
@@
Figure 4 shows the function fitting of step 2 on the ladder models, and the prediction of task accuracy for the target models using their actual task loss as the intermediate feature. For all tasks, all data points fit well with a shared sigmoidal function, with the average relative fitting error ranging from 0.4% to 2.6%. We get within 3% relative prediction error on MMLU, HellaSwag, PIQA, and CommonsenseQA. Overall, we get an average relative error of 3.7% for 7B-4T, and 3.5% for 13B-5T.